OpenStudy (anonymous):

WILL FAN AND MEDAL FOR HELP Which equation is represented by the table? x y (x, y) 0 –1 (0, –1) 1 4 (1, 4) 2 9 (2, 9) A. y = 3x + 3 B. y = 5x – 1 C. y = 3x + 3 D. y = 3x – 4

OpenStudy (anonymous):

its B -1 = 5(0) -1 4 = 5(1) - 1 9 = 5(2) -1

OpenStudy (anonymous):

thx and i was thinking A or B so thankyou